# Altermagnetic anomalous Hall effect emerging from electronic correlations

### v1

@@ Abstract
where $h(\mathbf{k})=(b_{\mathbf{k}},c_{\mathbf{k}},h_{z})$, $h_{z}=-\Delta M^{\text{AFM}}_{z}\sigma_{z}$, $\Delta=\frac{3}{2}J_{z}+\frac{3}{4}|J_{z}|$, $\mu$ is the chemical potential, $\epsilon=\frac{3}{8}|J_{z}|(n-1)$, $n=n_{a}=n_{b}$ is the electron density of an atom A or B, with $n_{l}=\langle\hat{n}_{l\uparrow}\rangle+\langle\hat{n}_{l\downarrow}\rangle$, $l=a,\,b$ being the sublattice index. The antiferromagnetic $M^{\text{AFM}}_{z}$ order parameter along the $z$ direction is expressed in terms of the on-site magnetization as $M^{\text{AFM}}_{z}=\frac{1}{2}\left(m^{a}_{z}-m^{b}_{z}\right)$, where $m^{l}_{z}=\frac{1}{2}\left(\langle\hat{n}_{l\uparrow}\rangle-\langle\hat{n}_{l\downarrow}\rangle\right)$. The hopping terms $a_{\mathbf{k}}$, $b_{\mathbf{k}}$ and $c_{\mathbf{k}}$ are given by $a_{\mathbf{k}}=2\lambda\sum_{i=1}^{3}\sin\left(\mathbf{k}\cdot\mathbf{a}_{i}\right)$, $b_{\mathbf{k}}=t\sum_{i=1}^{3}\cos(\mathbf{k}\cdot\mathbf{\delta}_{i})$, $c_{\mathbf{k}}=t\sum_{i=1}^{3}\sin(\mathbf{k}\cdot\mathbf{\delta}_{i})$, $\mathbf{\delta}_{i}$ and $\mathbf{a}_{i}$ are, respectively, the vectors connecting NN and NNN sites; $\mathbf{a}_{1}=\sqrt{3}a(-\frac{1}{2},\frac{\sqrt{3}}{2})$, $\mathbf{a}_{2}=\sqrt{3}a(-\frac{1}{2},-\frac{{\sqrt{3}}}{2})$, $\mathbf{a}_{3}=\sqrt{3}a(1,0)$, $\mathbf{\delta}_{1}=a(\frac{\sqrt{3}}{2},\frac{1}{2})$, $\mathbf{\delta}_{2}=a(-\frac{\sqrt{3}}{2},\frac{1}{2})$, $\mathbf{\delta}_{3}=a(0,-1)$, where $a$ is the distance between NN sites. $\tau_{0}$ ($\sigma_{0}$) is the $2\times 2$ identity matrix in the sublattice (spin) space.
The energies of the mean field Hamiltonian (Eq. 9) are given by
@@
$E^{\text{MF}}_{\alpha,\sigma}({\mathbf{k}})=\epsilon-\mu+a_{\mathbf{k}}\sigma+\alpha\sqrt{b^{2}_{\mathbf{k}}+c^{2}_{\mathbf{k}}+\left(\Delta M^{\text{AFM}}_{z}\right)^{2}},$
@@
where $\alpha=\pm$ ($\sigma=\pm$) is the band (spin) index. The spin-dependent mass term $h_{z}$ in Eq. (9) breaks TRS and a non-vanishing Berry curvature (BC) is expected away from half-filling. For a given spin projection $\sigma$, the BC of a band $\alpha$ can be expressed as [26, 27]
@@
$\Omega^{\alpha}_{z}(\sigma,\mathbf{k})=-\frac{\alpha}{2|\mathbf{h}({\mathbf{k}})|^{3}}\mathbf{h}(\mathbf{k})\cdot\left[\partial_{k_{x}}\mathbf{h}(\mathbf{k})\times\partial_{k_{y}}\mathbf{h}(\mathbf{k})\right]$
and the anomalous Hall (AH) conductivity [5] is given by the integral
@@
$f_{\alpha}(\mathbf{k},\mu)$ being the Fermi-Dirac distribution function  [28]$\sigma_{\rm Hall}$ can be computed using the values of the chemical potential and the AFM magnetization $M^{\text{AFM}}_{z}$ extracted from the mean field calculations. The results are depicted in Fig. 3 showing $\sigma_{\rm Hall}$ as function of the electronic density $n$.
The occurrence of a non-vanishing AH conductivity can be understood from the spin-valley dependence of the BC and the offset of the Dirac points induced by the modified Kane-Mele term $a_{\mathbf{k}}\sigma_{z}\tau_{0}$ (Eq. 9) which shifts oppositely the spin-split bands. To gain insights into these features, let us focus on the states $\mathbf{k}=\mu_{z}\mathbf{K}+\mathbf{q}$, ($\mathbf{q}\ll|\mathbf{K}|$) in the vicinity of the Dirac points $\mu_{z}\mathbf{K}$, where $\mu_{z}=\pm$ is the valley index. These states contribute mainly to the BC which reduces, around $\mu_{z}\mathbf{K}$, to
$\Omega^{\alpha}_{z}(S_{z},\mathbf{q})=\operatorname{sgn}\frac{v_{F}\Delta M^{\text{AFM}}_{z}}{(\hbar v_{F}q)^{2}+\left(\Delta M^{\text{AFM}}_{z}\right)^{2}},$ (10)
where $\operatorname{sgn}=-\alpha\mu_{z}\sigma$ is the sign of the BC and $\hbar v_{F}=\frac{3}{2}at$
Given the spin-dependent offset of the Dirac points, a spin-split sub-band $E^{\text{MF}}_{\alpha,\sigma}({\mathbf{k}})$ ($E^{\text{MF}}_{\alpha,-\sigma}({\mathbf{k}})$) of an occupied band $\alpha$, contributes to the BC by the states around the valley $\mu_{z}$ ($-\mu_{z}$), which results into a non-vanishing total BC, away from half filling.
The doping dependence of the AH conductivity (see Fig. 3) reflects the behavior of the secondary order parameter of Haldane type $N_{\rm AQH}$. At half-filling ($n=1$), $\sigma_{\rm Hall}$ vanishes as the bands below the Fermi level are fully occupied, resulting into a zero BC. By decreasing $n$ from half-filling, $\sigma_{\rm Hall}$ increases up to a maximum value, at a doping $n_{c}$, then it drops and vanishes below a critical doping value $n_{c0}$. The larger the coupling constant $J_{z}$, the smaller $n_{c0}$. For a fixed $J_{z}$, the increase of $\sigma_{\rm Hall}$ can be understood from the BC contribution which is enhanced as the area of the spin-polarized Fermi surface increases. The drop of $\sigma_{\rm Hall}$ is a consequence of the sharp decrease of $M^{\text{AFM}}_{z}$ below $n_{c}$, inducing a decrease of the BC The doping regime with a non-zero AH conductivity gets wider as $J_{z}$ increases, which results from the enhancement of the mass term $M^{\text{AFM}}_{z}$.
@@
Conclusions and Outlook — We have shown how electronic interactions induce altermagnetism in a non-centrosymmetric system by stabilizing a primary AFM ordering that hosts a secondary one that directly induces the altermagnetic anomalous Hall effect. Specifically, our quantum Monte Carlo calculation on interacting modified Kane-Mele lattice model reveal a finite temperature phase transition into an altermagnetically ordered state, whose physical features are captured by its effective continuum theory. The interacting model may in principle be implemented in cold atoms in optical lattices which have been proposed to realize altermagnetism [29]. Interestingly, recently an experimental realization of a non-centrosymmetric 3D altermagnet has been reported in GdAlSi, a collinear antiferromagnetic Weyl semimetal [30]. This suggests generalization of the lattice model to higher dimensions to determine how electronic correlations generate Berry curvature in altermagnetic semi-metals that may as well harbor 3D electronic topology.
@@ Acknowledgements.
We acknowledge fruitful discussions with Paul McClarty, Olena Gomonay, and Libor Šmejkal. We gratefully acknowledge the Gauss Centre for Supercomputing e.V. (www.gauss-centre.eu) for funding this project by providing computing time on the GCS Supercomputer SUPERMUC-NG at Leibniz Supercomputing Centre (www.lrz.de), (project number pn73xu) as well as the scientific support and HPC resources provided by the Erlangen National High Performance Computing Center (NHR@FAU) of the Friedrich-Alexander-Universität Erlangen-Nürnberg (FAU) under the NHR project b133ae. NHR funding is provided by federal and Bavarian state authorities. NHR@FAU hardware is partially funded by the German Research Foundation (DFG) – 440719683. FA, TS, JvdB and ICF thank the Würzburg-Dresden Cluster of Excellence on Complexity and Topology in Quantum Matter ct.qmat (EXC 2147, project-id 390858490). S. H. acknowledges the Institute for Theoretical Solid State Physics at IFW (Dresden) and the Max Planck Institute for the Physics of Complex Systems (Dresden) for kind hospitality and financial support.

### v2

@@ Abstract
where $h(\mathbf{k})=(b_{\mathbf{k}}{\sigma^{0}},c_{\mathbf{k}}{\sigma^{0}},h_{z})$, $h_{z}=-\Delta M^{\text{AFM}}_{z}\sigma^{z}$, $\Delta=\frac{3}{2}J_{z}+\frac{3}{4}|J_{z}|$, $\mu$ is the chemical potential, $\epsilon=\frac{3}{8}|J_{z}|(n-1)$, $n=n_{a}=n_{b}$ is the electron density on site A or B, with $n_{l}=\langle\hat{n}_{l\uparrow}\rangle+\langle\hat{n}_{l\downarrow}\rangle$, $l=a,\,b$ being the sublattice index. The antiferromagnetic $M^{\text{AFM}}_{z}$ order parameter along the $z$ direction is expressed in terms of the on-site magnetization as $M^{\text{AFM}}_{z}=\frac{1}{2}\left(m^{a}_{z}-m^{b}_{z}\right)$, where $m^{l}_{z}=\frac{1}{2}\left(\langle\hat{n}_{l\uparrow}\rangle-\langle\hat{n}_{l\downarrow}\rangle\right)$. The hopping terms $a_{\mathbf{k}}$, $b_{\mathbf{k}}$ and $c_{\mathbf{k}}$ are given by $a_{\mathbf{k}}={-}2\lambda\sum_{i=1}^{3}\sin\left(\mathbf{k}\cdot\mathbf{a}_{i}\right)$, $b_{\mathbf{k}}={-}t\sum_{i=1}^{3}\cos(\mathbf{k}\cdot\boldsymbol{\delta}_{i})$, $c_{\mathbf{k}}={-}t\sum_{i=1}^{3}\sin(\mathbf{k}\cdot\boldsymbol{\delta}_{i})$. $\boldsymbol{\delta}_{i}$ and $\mathbf{a}_{i}$ are, respectively, the vectors connecting NN and NNN sites; $\mathbf{a}_{1}=\sqrt{3}a(-\frac{1}{2},\frac{\sqrt{3}}{2})$, $\mathbf{a}_{2}=\sqrt{3}a(-\frac{1}{2},-\frac{{\sqrt{3}}}{2})$, $\mathbf{a}_{3}=\sqrt{3}a(1,0)$, $\boldsymbol{\delta}_{1}=a(\frac{\sqrt{3}}{2},\frac{1}{2})$, $\boldsymbol{\delta}_{2}=a(-\frac{\sqrt{3}}{2},\frac{1}{2})$, $\boldsymbol{\delta}_{3}=a(0,-1)$, where $a$ is the distance between NN sites. $\tau^{0}$ ($\sigma^{0}$) is the $2\times 2$ identity matrix in the sublattice (spin) space.
The spectrum of the mean field Hamiltonian (9) is given by
@@
$E^{\text{MF}}_{{\nu},\sigma}({\mathbf{k}})=\epsilon-\mu+a_{\mathbf{k}}\sigma+{\nu}\sqrt{b^{2}_{\mathbf{k}}+c^{2}_{\mathbf{k}}+\left(\Delta M^{\text{AFM}}_{z}\right)^{2}},$
@@
where ${\nu}=\pm$ ($\sigma=\pm$) is the band (spin) index. The spin-dependent mass term $h_{z}$ in Eq. (9) breaks TRS and a nonvanishing Berry curvature (BC) is expected away from half-filling. For a given spin projection $\sigma$, the BC of a band ${\nu}$ can be expressed as [29, 30]
@@
$\Omega^{{\nu}}_{z}(\sigma,\mathbf{k})=-\frac{{\nu}}{2|\mathbf{h}({\mathbf{k}})|^{3}}\mathbf{h}(\mathbf{k})\cdot\left[\partial_{k_{x}}\mathbf{h}(\mathbf{k})\times\partial_{k_{y}}\mathbf{h}(\mathbf{k})\right]$
and the AH conductivity [7] is given by the integral
@@
where $h(\mathbf{k})=(b_{\mathbf{k}},c_{\mathbf{k}},-\Delta M^{\text{AFM}}_{z}\sigma)$ and $f_{{\nu}}(\mathbf{k},\mu)$ is the Fermi-Dirac distribution function  [31].
$\sigma_{\rm Hall}$ can be computed using the values of the chemical potential and the AFM magnetization $M^{\text{AFM}}_{z}$ extracted from the mean field calculations. The results are depicted in Fig. 3 showing $\sigma_{\rm Hall}$ as function of the electronic density $n$.
The occurrence of a nonvanishing AH conductivity can be understood from the spin-valley dependence of the BC and the offset of the Dirac points induced by the modified KM term $a_{\mathbf{k}}\sigma^{z}\tau^{0}$ of Eq. (9), which shifts oppositely the spin-split bands. To gain further insight into these features, let us focus on the states $\mathbf{k}=\mu_{z}\mathbf{K}+\mathbf{q}$, ($\mathbf{q}\ll|\mathbf{K}|$) in the vicinity of the Dirac points $\mu_{z}\mathbf{K}$, where $\mu_{z}=\pm$ is the valley index. These states contribute strongly to the BC, which reduces, around $\mu_{z}\mathbf{K}$, to
$\Omega^{{\nu}}_{z}(\sigma,\mu_{z},\mathbf{q})=\operatorname{sgn}\frac{(\hbar v_{F})^{2}\Delta M^{\text{AFM}}_{z}}{2\left[(\hbar v_{F}\mathbf{q})^{2}+\left(\Delta M^{\text{AFM}}_{z}\right)^{2}\right]^{3/2}},$ (10)
where $\operatorname{sgn}=-{\nu}{\mu_{z}}\sigma$ is the sign of the BC and $\hbar v_{F}=\frac{3}{2}at$. Given the spin-dependent offset of the Dirac points, a spin-split subband $E^{\text{MF}}_{{\nu},\sigma}({\mathbf{k}})$ [$E^{\text{MF}}_{{\nu},-\sigma}({\mathbf{k}})$] of an occupied band ${\nu}$, contributes to the BC by the states around the valley $\mu^{z}$ ($-\mu^{z}$). Thus, the two contributing spin-dependent BCs have the same sign, resulting in a nonvanishing total BC away from half-filling. Data are shown in the Supplemental Material [32]. We note that in the continuum limit, our modified KM term, $a_{\mathbf{k}}\tau^{0}\sigma^{z}$, indeed corresponds to a valley-Zeeman spin-orbit coupling, akin to that considered in transition metal dichalcogenides [33, 34]. Both systems exhibit effects that induce spin-split band structures and enable spin-valley locking, but their underlying origins differ, as detailed in the Supplemental Material [32].
The doping dependence of the AH conductivity (see Fig. 3) reflects the behavior of the secondary order parameter of Haldane type $M_{\rm H}$. At half-filling ($n=1$), $\sigma_{\rm Hall}$ vanishes as the bands below the Fermi level are fully occupied, resulting in zero total BC. Decreasing $n$ from half-filling, $|\sigma_{\rm Hall}|$ increases up to a maximum value, at a doping $n_{c}$, where it drops and vanishes below a critical doping value $n_{c0}$. Larger coupling constants $J_{z}$ give smaller $n_{c0}$. For a fixed $J_{z}$, the increase of $|\sigma_{\rm Hall}|$ can be understood from the BC contribution which is enhanced as the area of the spin-polarized Fermi surface increases (see the Supplemental Material [32]). The drop of $|\sigma_{\rm Hall}|$ is a consequence of the sharp decrease of $M^{\text{AFM}}_{z}$ below $n_{c}$, inducing a decrease of the BC. The doping regime with a nonzero AH conductivity gets wider as $J_{z}$ increases, which results from the enhancement of the mass term $M^{\text{AFM}}_{z}$.
@@
Conclusions and outlook — We have shown how electronic interactions induce altermagnetism in a noncentrosymmetric system by stabilizing a primary AFM ordering that hosts a secondary one that directly induces the altermagnetic anomalous Hall effect. Specifically, our quantum Monte Carlo calculations on the interacting modified Kane-Mele lattice model reveal a finite temperature phase transition into an altermagnetically ordered state, whose physical features are captured by its effective continuum theory. The interacting model may in principle be implemented in cold atoms in optical lattices that have been proposed to realize altermagnetism [35]. Time-modulated optical lattices of ultracold atoms allow for complex NNN tunneling, as demonstrated for the Haldane model [36], while Hubbard models with AFM order have been realized in ultracold fermions [37, 38]. This suggests potential for combining these techniques, incorporating both complex NNN tunneling and Hubbard-type models with AFM order. Interestingly, recently an experimental realization of a non-centrosymmetric 3D altermagnet has been reported in GdAlSi, a collinear antiferromagnetic Weyl semimetal [39]. This suggests generalization of the lattice model to higher dimensions to determine how electronic correlations generate Berry curvature in altermagnetic semimetals that may as well harbor 3D electronic topology.
@@ Acknowledgements.
Acknowledgments — We acknowledge fruitful discussions with Paul McClarty, Olena Gomonay, Libor Šmejkal, and Tomáš Jungwirth. We gratefully acknowledge the Gauss Centre for Supercomputing e.V. for funding this project by providing computing time on the GCS Supercomputer SUPERMUC-NG at Leibniz Supercomputing Centre, (project No. pn73xu) as well as the scientific support and HPC resources provided by the Erlangen National High Performance Computing Center (NHR@FAU) of the Friedrich-Alexander-Universität Erlangen-Nürnberg (FAU) under the NHR project b133ae. NHR funding is provided by federal and Bavarian state authorities. NHR@FAU hardware is partially funded by the German Research Foundation (DFG) — 440719683. FA, TS, JvdB and ICF thank the Würzburg-Dresden Cluster of Excellence on Complexity and Topology in Quantum Matter ct.qmat (EXC 2147, project-id 390858490). S. H. acknowledges the Institute for Theoretical Solid State Physics at IFW (Dresden) and the Max Planck Institute for the Physics of Complex Systems (Dresden) for kind hospitality and financial support.